### v4

# On Diversified Preferences of Large Language Model Alignment

Abstract
Aligning large language models (LLMs) with human preferences has been recognized as the key to improving LLMs’ interaction quality. However, in this pluralistic world, human preferences can be diversified due to annotators’ different tastes, which hinders the effectiveness of LLM alignment methods. This paper presents the first quantitative analysis of commonly used human feedback datasets to investigate the impact of diversified preferences on reward modeling. Our analysis reveals a correlation between the calibration performance of reward models (RMs) and the alignment performance of LLMs. We find that diversified preference data negatively affect the calibration performance of RMs on human-shared preferences, such as Harmless&Helpful, thereby impairing the alignment performance of LLMs. To address the ineffectiveness, we propose a novel Multi-Objective Reward learning method (MORE) to enhance the calibration performance of RMs on shared preferences. We validate our findings by experiments on three models and five human preference datasets. Our method significantly improves the prediction calibration of RMs, leading to better alignment of the Alpaca-7B model with Harmless&Helpful preferences. Furthermore, the connection between reward calibration and preference alignment performance suggests that calibration error can be adopted as a key metric for evaluating RMs. The open-source code and data are available at https://github.com/dunzeng/MORE.
On Diversified Preferences of Large Language Model Alignment
Dun Zeng1,333Work done during an internship at Tencent AI Lab,111Equal contribution. YongDai1,111Equal contribution. Pengyu Cheng1,111Equal contribution. Longyue Wang1 Tianhao Hu1 Wanshun Chen1 Nan Du1 Zenglin Xu2,222Corresponding author: zenglin@gmail.com 1Tencent AI Lab, Shenzhen, China 2Peng Cheng Lab, Shenzhen, China
1 Introduction
Large language models (LLMs), such as ChatGPT OpenAI (2023) and LLaMa (Touvron et al., 2023a, b), have significantly accelerated the development process toward artificial general intelligence (AGI). Among the key factors for such great achievement, the alignment technique, which finetunes LLMs with human feedback (Christiano et al., 2017), has played an essential role in training LLMs’ responses to follow human values (e.g., helpfulness and harmlessness) (Askell et al., 2021). Among the LLM alignment algorithms, reinforcement learning from human feedback (RLHF) (Ouyang et al., 2022) has become the mainstream solution, which first learns a reward model (RM) representing human preferences and then updates LLMs via the proximal policy optimization (PPO) (Schulman et al., 2017) toward generating responses with higher RM scores. Alternative alignment methods also have been sequentially proposed for better computational complexity and training instability, such as RAFT Dong et al. (2023b), DPO Rafailov et al. (2023), RRHF Yuan et al. (2023), and APO Cheng et al. (2023b).
The performance of these alignment methods highly depends on the quality of human preference data , where is the input query to the LLM, and response is preferred to response under the human annotation Ouyang et al. (2022). Ideally, the preference datasets should uniformly be helpful, harmless, benevolent, and unbiased to guide the LLM alignment. However, in real-world scenarios, individuals can have diversified preferences on the same topic based on their different experiences, educational backgrounds, religions, and cultures (Leonardelli et al., 2021). Even for the same person, his or her expected model answer to a particular question can vary depending on different scenarios (Cheng et al., 2023a). The annotation disagreement, which is caused by different annotators or the same annotator in different scenarios Bai et al. (2022), will significantly hinder the effectiveness of alignment methods Davani et al. (2022); Wan et al. (2023).
To identify the diversified preferences quantitatively, we select five commonly used human feedback datasets, train an RM on each, and then test the performance on the other sets (details in Section 3). We plot the observation results in Figure 1. We observe that training RM on a single preference data source may cause inconsistent reward distribution shifts (middle plot), result in diverse reward values (right plot), and compromise the performance of other sets (left plot). The result indicates that different human preference datasets have different preference distributions (Cheng et al., 2023a). Hence, a more comprehensive understanding of the impact of diversified human preference datasets on the reward model becomes crucial, yet it has not received adequate attention and remained unexplored in the LLM alignment domain.
In our exploration, we found the over-rewarding phenomenon, that is, the vanilla RMs tend to output extreme rewards on samples, which damages the RMs and LLM alignment. To enhance the efficiency of leveraging the diversified preference datasets, inspired by multi-objective optimization methods Sener and Koltun (2018); Zeng et al. (2023), we regard RMs as a shared reward additionally with a customized reward drift. The shared reward represents the shared preferences across datasets (or general human preferences) and the reward drift contains individual or domain-specific preference information (Cheng et al., 2023a). Then, we introduce a Multi-Objective Reward training scheme (MORE) to capture the shared (general) preference information, which adopts a novel reweight techniques to minimize the mean gradient of enlarging reward drifts. With MORE, RMs can capture a broader range of preferences and mitigate the impact of reward drifts. The main contributions of this paper are:
This is the first work to demonstrate the positive correlation between the calibration performance of RMs and the alignment performance of LLMs. Moreover, RM learning on diversified preferences typically induces high calibration errors, indicating unreliable rewards. The unreliable rewards come from a over-rewarding phenomenon, denoting vanilla RMs output extreme rewards inducing harmful reward drifts. Hence, it negatively impacts the performance of LLM alignment.
We induce a simple and effective Multi-Objective Reward (MORE) training scheme to alleviate the over-rewarding phenomenon. MORE makes self-adaption to the RM learning gradient to mitigate the reward drifts. MORE effectively enhances the calibration performance of RMs, especially on shared preferences across diversified preference datasets.
We verified our findings with Pythia-1.4B, Pythia-2.8B (Biderman et al., 2023) and LLaMa2-7B (Touvron et al., 2023b) on five widely recognized and diverse preference datasets. Through empirical analysis, we established that MORE significantly minimizes reward drift and achieves low Expected Calibration Error (ECE) values. Additionally, by applying reject sampling to Alpaca-7B (Taori et al., 2023) with the RMs generated, we aligned the models with Helpful&Harmless preferences, thereby affirming the critical role of ECE in the evaluation of Reward Models.
2 Background
Large language Model Alignment
Parameterized by , a reward model (RM) is a mapping , which provides a real-valued reward score evaluating a textual response corresponding to an input prompt . Given a sample from a preference dataset , is expected to provide a preference score with , representing the response is preferred. Following the Bradley-Terry model (David, 1963), the RM learning objective on the preference dataset is defined as:
where we use to denote reward difference for simplifying notation in this paper and is the Sigmoid function. With a well-learned reward , LLM alignment optimizes the generation policy by maximizing the expected reward value:
where is the KL divergence regularizer between current policy and a reference , preventing the optimization from instability and degeneration. The typical solution to the preference optimization in equation 3 is reinforcement learning (RLHF) (Ouyang et al., 2022), especially with the proximal policy optimization (PPO) algorithms (Schulman et al., 2017). However, RLHF has been recognized as practically suffering from implementation complexity and training instability. To avoid the RL schedule during alignment, reject sampling methods (Liu et al., 2023) directly conduct supervised fine-tuning on to further simplify the human preference alignment process. The rejection sampling optimization (RJS) loss can be written as
where is the sampled response with the highest reward score.
Calibration Error
Calibration error is an effective method to estimate the confidence of a model’s outputs (Guo et al., 2017). We divide the confidence interval with finite samples into bins with equal length (). Then, we place model predictions into these bins according to their prediction confidence. Let be the set of indices of samples that fall into the internal . We calculate the corresponding accuracy and average confidence of each bin as follows:
where are the prediction results, and is the ground-truth of the -th sample. is the indicator function which produces 1 if otherwise 0. is the prediction confidence of the -th sample. In the context of reward modeling, the prediction confidence in (1). For a set of samples, we can compute the Expected Calibration Error as follows:
We set for measuring calibration performance in this paper.
Numerous studies have focused on improving the calibration performance of statistical machine-learning systems (DeGroot and Fienberg, 1983; Palmer et al., 2008; Yang and Thompson, 2010). Furthermore, the calibration error of neural networks provides additional information for users to determine whether to trust the model’s predictions, especially for modern neural networks that are more challenging to interpret (Guo et al., 2017; Zhu et al., 2023). In the field of natural language processing, studies have revealed a positive relationship between calibration performance and the reduction of hallucination (Xiao and Wang, 2021; Tian et al., 2019), and the evaluation of pre-trained language models (Kadavath et al., 2022; Tian et al., 2023). The calibration error has demonstrated its ability to evaluate the performance of language models. In this paper, we first employ the calibration error to evaluate the RMs. Subsequently, we investigate the implicit connection between RMs and LLM alignment under diversified preferences.
3 Empirical Study of Diversified Preferences
We start with an empirical analysis of diversified preferences in reward modeling on multiple sources , where each data source contains the preference comparison pairs from different tasks (Dong et al., 2023a), domains (Cheng et al., 2023a), or individuals Bai et al. (2022). In this paper, we selected Summarize (Stiennon et al., 2020), Webgpt (Nakano et al., 2021a), Helpful&Harmless (Bai et al., 2022), and OASST1 (Köpf et al., 2023) as the different preference sources to empirical analysis the phenomena of diversified preferences. We use Pythia-1.4B (Biderman et al., 2023) as the RM base, and finetuned RMs with comparisons from each source. The experiment setup aligns with Section 5.
The reward distributions across various RMs exhibit diversity when applied to the same dataset. We analyze and present the variation in rewards (defined as the difference in reward values assigned by an RM to the winning and losing samples) offered by these RMs, as illustrated in Figure 1 (additional results in Figure 8 and LABEL:fig:additional-results in Appendix). Compared with the results of raw model , we observe that training on different datasets results in diverse reward values (right plot) and distribution shift (middle plot). Specifically, the reward value distribution of shifts from the in a certain degree. While the reward value distributions of , , and shifts to the a different direction. Moreover, despite the distribution of , , and are similar, the mean-variance of their reward values are quite different.
Furthermore, when considering the accuracy gains illustrated in Figure 1 (left plot), the observed shift in reward distribution indicates that the learned reward values from preference datasets are diversified. To effectively capture the shared reward values across these diversified preferences, it becomes necessary to formulate a new problem approach for reward modeling on diverse preference datasets.
4 Multi-Objective Reward Learning
In this section, we propose our reward modeling on diversified preference datasets, highlighting the implicit reward drift during the reward learning process and its negative impacts. Then, we present the MORE training schemes to mitigate the reward drifts as a feasible solution. To maintain the integrity of our paper, we leave our quantitative analyses of reward modeling on diversified preferences in the next section.
4.1 Preference Diversity as Reward Drift
We denote as the shared reward function, which (ideally) provides reward values reflecting the general values among people (or shared preference information across datasets in practice). As the collected human-feedback datasets are limited and implicitly biased, training an RM on a limited preference dataset can be viewed as drifting from an optimal reward. We can form a reward model with reward drift in a data level:
where , and is the reward drift learned by RM . Then, we investigate the vanilla ranking loss for reward modeling. Substituting reward function in (1) with the drifted form (4), we have
Hence, updating the RM to minimize the rank loss will enlarge the reward differences (input of the Sigmoid function). Simultaneously, the reward drift is also enlarged, causing over-rewarding.
4.2 Reward Modeling on Diversified Data
Letting be the RM trained on mixed diverse datasets , the can be viewed as a multi-task learner with shared parameters (Sener and Koltun, 2018). Then, the reward value provided by can be decomposed into voting format weighted by an implicit :
where the shared reward is the same with arbitrary , and is the reward drift. We interpret that the is provided by subset of parameters , representing the preferences from the -th dataset . This reward value decomposition naturally holds in the model output level, despite the non-linear nature of neural networks.
Moreover, our formulation aligns with multi-task learning (Crawshaw, 2020) and multi-objective learning (Guardieiro, 2023) problems. For example, the can be implemented as an ensemble model, where is the base models. Therefore, it is natural to adjust the weight in an ensemble manner (Coste et al., 2023; Jang et al., 2023; Touvron et al., 2023a; Eisenstein et al., 2023) to mitigate the reward drift such that . Compared with average rewards from multiple RMs Jang et al. (2023); Eisenstein et al. (2023), we focus on training a single RM that learns the shared preference. We propose to reduce the model update on reward drift during RM training via linear scalarization Barrett and Narayanan (2008). Moreover, we provide further discussion on related manners in Section 7.
4.3 Training Scheme: MORE
MORE loss function
Our analyses suggest finding proper weights for mitigating reward drifts. Then, we propose training RMs to capture the shared preference across multiple datasets with the following objective:
where . Compared with vanilla ranking loss in (1), the above loss additionally focuses on the combination relation across preferences. The linear combination of loss functions is commonly adopted in deep learning methods to balance the interaction of different modules (Zhang et al., 2023; Kurin et al., 2022). Analogously, we treat each preference as an individual module and balance them wisely. Moreover, this formulation also covers several typical training cases. For example, directly mixing diverse preference datasets and training a RM implicitly induces (McMahan et al., 2017; Ramé et al., 2024). Therefore, if the number of data samples from a single preference is greatly larger than other preferences, the RM is likely to drift to the preference with more samples. Excluding data quantity, the weight is also decided by the quality of data samples in the training process (Katharopoulos and Fleuret, 2018; Zhou and Wu, 2023). Neural network training typically provides a larger gradient for harder samples (Katharopoulos and Fleuret, 2018), therefore, leaning the RMs preferences drift to these hard samples. In practice, the quantity and quality variance in diversified datasets may require more hyper-parameter searching (Guo et al., 2024) or data composition efforts (Dong et al., 2023a) in the vanilla finetuning process.
What is MORE doing?
We suggest training a better RM via self-adaption training weights for better data efficiency. The MORE loss minimizes the ranking loss by solving a reward drift mitigation task, applying a batch-wise reweighting method. Let batch data be the sampled batch data from diverse datasets. Furthermore, is the subset of batch data from the -th preference dataset. We have the gradient
where we adjust to minimize the partial gradient of enlarging reward drifts. The mitigation task in (8) can be efficiently solved by the Frank-Wolfe solver (Jaggi, 2013; Sener and Koltun, 2018; Zhou et al., 2022b; Zeng et al., 2023). We provide the details of our efficient implementation in the Appendix B. Furthermore, shares the same magnitude of vanilla loss function in expectation over the whole training dataset, as justified in Appendix B.
Outline
The MORE only requires simple modification on batch data sampling and batch-wise reweighting. We depict the pipeline in Figure 2. MORE consists of THREE main steps as: 1) Sample a diverse batch data , and input the batch data forward the RM and obtain the hidden states , which is the inputs of the reward head . 2) Compute the gradient of reward head with data . 3) Compute the weights by Frank-Wolfe solver. Finally, we substitute the loss weights in (7) as the final loss for the optimizer to conduct backward and model updating. This procedure prevents the RM from enlarging implicit reward drifts.
5 Experiments on Reward Modeling
In this section, we present our experiments and quantitative analyses on reward modeling.
Datasets & models
We use open-sourced human preference alignment datasets, including Helpful&Harmless (Bai et al., 2022), OASST1 (Köpf et al., 2023), Webgpt (Nakano et al., 2021a), and Summarize (Stiennon et al., 2020). We provide the statistics of the datasets and data composition in Appendix 3. Despite these datasets being released to human preference alignment, our study highlights the preference diversity across the datasets and its impacts on training RMs. We train Pythia-1.4B, Pythia-2.8B (Biderman et al., 2023) and LLaMa2-7B (Touvron et al., 2023b) as the LM base for RM training. We use the last token embedding of the output hidden states as the pooled hidden representation, then add one linear layer (RM head) with the scale-value output on it to predict reward scores. We present the details of the training setup in Appendix C.
Baselines
We compare our method with conventional fine-tuning strategies for training language models, specifically mixing the preference data samples. We refer to the training scheme as MultiTask training Dong et al. (2023a). The MultiTask training scheme randomly samples data from hybrid preference datasets. Additionally, we compare with the Top performance of RMs trained on each preference dataset. We highlight that the Top performance indicates the ideal ensemble-RM, i.e., each sample obtains its reward from the corresponding best RM. Then, we naively Average the reward values from Top RMs provide on the same samples to denote a naive ensemble-RM. In all, we mark the baseline rewards as RM, RM and RM respectively.
Evaluation metric
We use the preference accuracy on test datasets for each domain. If an RM outputs for a test sample , we denote it as a correct prediction. The preference accuracy is then computed as the proportion of correct predictions within all testing response pairs. However, preference accuracy only provides pairwise comparisons of responses and does not reflect the degree of preference for each response. Following Bai et al. (2022); Cheng et al. (2023b), we examine the probability calibration to test if the learned RMs accurately represent the human preference distribution. This is measured by the Expected Calibration Error (Naeini et al., 2015; Zhu et al., 2023).
5.1 Reward Modeling on Diversified Preference Datasets
We provide the reward modeling results on mixed diversified datasets in Figure 3 and Figure 4. The detailed information is in Table 2 of the Appendix.
The reward accuracy does not drop significantly on mixed diversified preferences.
Increasing the size of LLMs, reward model training on mixed diversified preference datasets can maintain reward accuracy. For instance, when Pythia-1.4B is used as the RM base model, the reward accuracy is lower compared to the Top accuracy achieved through single preference training on all preferences. Then, when LLaMa2-7B is used as the base model, the reward accuracy on the Oasst1, Webgpt, and Summarise test sets surpasses the top accuracy achieved through single training. Additionally, the degradation of reward accuracy on the Helpful and Harmless datasets is mitigated. Therefore, the performance of RMs typically is proportional to the size of base models (Gao et al., 2023). Moreover, we find the accuracy of RM is poor, revealing the preference conflicts across RM.
Reward modeling on mixed diversified preferences affects calibration performance
Noting the reward accuracy only provides comparisons of responses (Zhu et al., 2023), we emphasize the ECE performance reflects the degree of preference for responses in Figure 4. Compared RM with RM, reward modeling on mixing the diversified preference datasets typically degenerates calibration performance on all preferences. Especially, the reward accuracy of RM and RM are comparable but the calibration performances are very different. The LLMs can maintain high accuracy on all preferences due to their large capacity, however, the reward distribution is affected by mixed diversified preferences. These findings reveal that reward accuracy is insufficient to verify the ability of RMs and suggest evaluation of RMs via ECE. We will further justify the point in the alignment experiments.
Reward Model Perplexity (PPL) GPT4 Evaluation (%) Base Model Scheme Acc(%) ECE Helpful Harmless Win Tie Lose - - - - 15.48 12.71 - - - Pythia-1.4B MultiTask 64.79 0.0177 15.30 8.22 44 22 34 MORE 64.32 0.0109 12.68 8.42 Pythia-2.8B MultiTask 66.61 0.0145 16.76 8.42 45 21 34 MORE 65.87 0.0078 13.14 10.29 LLaMa2-7B MultiTask 72.40 0.0284 16.93 8.69 45 23 32 MORE 72.32 0.0143 11.97 9.96 Table 1: The RJS alignment performance with different RMs. The first line is the performance of Alpaca base model. The results show that ECE further reflects the ability of RMs when the reward accuracy is close.
Figure 7: The correlation between ECE of RMs and RJS alignment performance for the Alpaca model.
MORE implements significant calibration performance improvement
The RM preserves a significantly lower ECE than RM, indicating that RM provides more accurate reward values. Moreover, RM implements significantly lower ECE than RM on Helpful&Harmless preferences. This is because Helpful&Harmless preference is shared by these datasets and MORE accurately captures shared preferences across them. Therefore, MORE implements lower calibration errors on shared Helpful&Harmless preference and slightly loses its calibration performance on the other three preference datasets. This calibration performance gap between RM and RM on the other three diversified preferences further reflects the preference diversity.
5.2 Analyses on RMs of H&H Preferences
To clarify the improvement of MORE, we provide analyses on Helpful&Harmless (H&H) datasets, which is an important human preference alignment objective for LLMs in recent works (Ouyang et al., 2022; Touvron et al., 2023b). Concretely, we focus on the statistics of the reward difference (i.e., ). We count the reward differences of RMs on H&H test datasets in Figure 6.
MORE mitigates over-rewarding phenomenon
In Figure 6, we observe the RM outputs large absolute reward differences on testing samples. On the contrary, the RM provides lower absolute reward differences on testing samples, compared with baseline training schemes. Moreover, RMs tend to provide extreme rewards to some samples. We count these extreme reward values as outliners in Appendix, Table 6. This phenomenon aligns with our methodology in (6), that is, MORE mitigated the reward drifting during training. Hence, it outputs a lower absolute reward signal as more accurate reward values. These findings reveal the phenomenon of over-rewarding in RMs, where vanilla RMs tend to assign large reward values to samples. This phenomenon demonstrates problem modeling (5). Importantly, the over-rewarding in RM may not break the reward accuracy shown in Figure 3, however, it induces unsatisfied calibration performance. MORE maintains the reward accuracy of RMs, alleviates the over-rewarding effects on reward modeling, and trains better RMs.
MORE achieves better calibration using more diversified preferences
The MORE can benefit from diversified preference information by (8), which suggests increasing the number of diversified preferences can better mitigate reward drifts. We change the number of mixed preference datasets from 2 to 5 to verify our insights, as shown in Figure 6. In detail, we start from mixed Helpful&Harmless datasets (K=2) and then add Oasst1, Webgpt, Summarise datasets. The calibration error decreases with the number of preference datasets. It proves that MORE can utilize the preferences information to enhance the performance of the reward model on shared preferences and surprisingly outperforms RM.
6 Experiments on LLM Alignment
In this section, we use the previously obtained RMs for LLM alignment experiments on Alpaca (Taori et al., 2023), which is an instruction-tuned LLaMA-7B model (Touvron et al., 2023a). We use Reject Sampling (RJS) (Touvron et al., 2023b; Liu et al., 2023) as the alignment algorithms, where we sample 4 responses from Alpaca with queries from H&H trainsets. Our experiment mainly justifies the correlation between the calibration performance of RMs and LLM alignment performance.
RM works better than RM for RJS aligning H&H with lower ECEs
We finetune Alpaca with the most preferred samples scored by previously obtained RMs to align the human preference of H&H, following RJS loss (3). We show the alignment performance in Table 1, where we use the same GPT4 evaluation prompts with DPO (Zhou et al., 2023) shown in Appendix D. RM works better for RJS tasks. Noting that RM and RM implements comparable reward accuracy on H&H, while the calibration performance are significantly different. Therefore, the alignment performance is additionally related to the calibration performance of the RMs.
ECE of RMs is positively correlated with alignment performance
We finetune the Alpaca model on the good response from H&H training datasets, and the finetuned model is marked by Alpaca-SFT. Then, we conduct the RJS alignment experiments with LLaMa2-7B RMs from Figure 6. We compare each alignment result of Alpaca-RJS models with the same Alpaca-STF model via GPT evaluation, shown in Figure 7. The results show that the RMs with lower ECE values work better for RJS alignments, emphasizing the importance of calibration evaluation.
7 Additional Discussions
Connections with data composition and ensemble-RM studies
Dong et al. (2023a) have empirically shown that the LLM ability can be improved by adjusting the mixed training data ratio from different sources. However, the mixed proportion can be hard to search in practice. Besides, other studies have shown that direct ensemble RMs (Eisenstein et al., 2023) or merging RMs’ parameters (Jang et al., 2023; Ramé et al., 2024) during training could also improve the ability of RMs. In practice, these approaches induce a large system burden for storing/training multiple RMs, especially since the RMs can be extremely large. In comparison, this paper focuses on training single RM on diversified datasets.
Suggestions for reward model training
This paper reveals two main suggestions for future reward model training works. First, Evaluate RMs with reward accuracy and calibration error. Reward accuracy is insufficient to evaluate the ability of RMs due to model capacity and data quality. Our work suggests the community additionally focuses on the calibration performance of RMs. Besides, Increasing the diversity of preference data samples can ensure the robustness of the reward modeling process. Due to the preference information being typically noisy, learning reward information from mixed diversified datasets can be beneficial.
Applications
The MORE can enhance preference modeling pre-trained (PMP) paradigm (Askell et al., 2021) as it captures the shared preference information. This facilitates its use in federated learning scenarios (McMahan et al., 2017), where the data distributions are highly heterogeneous across participants. Moreover, the RM can be easily finetuned to specific preferences (Cheng et al., 2023a). This flexibility allows for the adaptation of our approach to various applications.
Extension to RM-free alignment methods
RM-free alignment methods (Rafailov et al., 2023; Azar et al., 2023) are derived based on an implicit reward model. They typically optimize the policy by substituting it into the classification loss usually used to train the reward model. The relation of calibration performance of implicit reward and the alignment performance in the RM-free methods is unexplored. Besides, learning shared preferences from mixed diverse preference datasets can be extended to RM-free paradigms. For example, we can re-weight the partial reward loss of the RM-free alignment methods, especially DPO (Rafailov et al., 2023; Zhou et al., 2023). We will explore this in future work.
8 Limitations
We only conducted experiments using the conventional RJS algorithm in LLM alignment tasks. As a reward modeling algorithm that captures shared preference information, MORE depends on the quality of the applied data. Therefore, the correlation of ECE of RMs and LLM alignment performance in other alignment algorithms requires further exploration. Besides, the training datasets we used contain violence, abuse, and biased content that can be upsetting or offensive to particular groups of people.
Appendix A Related Work
RLHF has become the mainstream approach to align language models towards helpfulness, and harmlessness (Leike et al., 2018; Nakano et al., 2021b; Ouyang et al., 2022; Bai et al., 2022). They all utilize an RM to align machine learning systems with human performance, which directly decides the performance of preference alignment. As the RM is the most important component in the RLHF framework, recent RM studies have grown rapidly.
Reward Modeling in human preference alignment
The original goal of RM is to provide a scalar score to a model response and indicate the quality in (2), especially helpfulness and harmlessness. Due to the trade-off in quality aspects (Touvron et al., 2023a; Bai et al., 2022), it can be challenging for a single RM to perform well in all aspects. Our work related to previous works handling multiple rewards and potential disagreement in preferences. For instance, LLaMa-2 (Touvron et al., 2023a) utilizes two separate RMs, one optimized for helpfulness and another for harmlessness. They mitigate the magnitude bias of the reward scalar with a margin loss, which provides a large margin for pairs with distinct responses, and a smaller one for those with similar responses. Multiple RMs can be utilized as majority voting or averaging (Jaques et al., 2020; Jang et al., 2023) in the PPO (Schulman et al., 2017). Wang et al. (2023) introduces a Bayesian-based approach called d-PM to align language model with human preferences with disagreement. Cheng et al. (2023a) proposes to train a customized RM from the general RM to avoid disagreement from different preference domains. Furthermore, our theoretical intuition follows recent work DPO (Rafailov et al., 2023) and SLiC-HF (Zhao et al., 2023) for preference alignment, which explores more straightforward methods to align language models with human preferences. Beyond the methodology, they have shown the RLHF framework is working as likelihood calibration tasks (Deng et al., 2020; Wang et al., 2023; Azar et al., 2023), which proves that the reward values provided by the RM are also important.
Domain Generalization
Machine learning methods suffer from performance degeneration when the source domain data and the target domain data follow different distributions, which has been recognized as the domain shift problem (Pan and Yang, 2009; Csurka, 2017; Wang et al., 2021). To address this problem, domain generalization is proposed to minimize the domain shift across domains. In this direction, existing methods aim to learn the domain invariant representation to reduce the discrepancy between representations of multiple source domains (Zhou et al., 2022a). We derive the concept of reward shift from domain shift. Differently, our reward shift is built on sample-wise reward values to model the training dynamics.
Multi-objective Optimization
Multi-objective Optimization (MOO) (Gunantara, 2018) is a branch of methods addressing learning problems involving multiple conflicting objectives. In real-world scenarios, it commonly encounters situations where multiple objectives need to be considered simultaneously, often with trade-offs between them. In the practice of machine learning, most MOO methods Sener and Koltun (2018); Zeng et al. (2023) apply linear scalarization Barrett and Narayanan (2008) to merge multiple objectives into one, and then automatically adjust the objective coefficients to balance the conflicts among different tasks.
Appendix B Detailed Discussions about MORE
Batch-wise reweighting
We use adaptive weighting methods to reduce the reward drift across preferences and adjust the reward modeling process in the data batch-wise. The mitigation task in (8) can be efficiently solved by the Frank-Wolfe solver (Jaggi, 2013; Sener and Koltun, 2018; Zhou et al., 2022b; Zeng et al., 2023). However, the computing cost of solving it is proportional to the size of parameters . Since the size of is in the billions, we only utilize gradients on the reward head from each preference to avoid expensive computation cost. In detail, we obtain the hidden states before the reward head and compute the gradient of the reward head solely with data . Collecting the reward head gradient from diversified preferences, the is computed by:
In this paper, we only utilize the gradient information on the reward head (simple linear layer). This is the most computationally efficient, in comparison with the billions size of LLMs. Moreover, there is a trade-off between gradient information utility and computation efficiency depending on the size of the utilized gradient (Sener and Koltun, 2018).
Decomposition of ranking loss
Using the properties of the sigmoid function and , we present the detailed decomposing of vanilla ranking loss gradients:
where we use the definition of reward drift in (4). Next, we decompose the second term of reward drifts:
where we induce the preference source of data samples in the last equation. Vanilla rank loss regards the importance of data samples as equal. Then, let us observe the gradient of MORE loss:
In comparison, the gradient replaces the coefficients with adjustable variable . Therefore, the vanilla ranking loss is a special case of MORE loss.
Appendix C Experiment Details
| Training | Testing Dataset (Acc %) | Metrics | |||||||
| Base Model | Dataset | Method | Helpful | Harmless | Oasst1 | Webgpt | Summ. | Avg. | ECE |
| Pythia-1.4B | - | Raw | 52.38 | 50.69 | 51.25 | 48.47 | 51.06 | 50.77 | 0.1281 |
| Single | Top | 67.81 | 69.07 | 62.43 | 65.70 | 62.56 | 65.51 | 0.0362 | |
| ALL | Averaging | 55.73 | 51.81 | 57.68 | 53.60 | 55.50 | 54.86 | 0.0543 | |
| ALL | MultiTask | 65.00 | 64.57 | 60.13 | 66.00 | 57.49 | 62.38 | 0.0541 | |
| ALL | MORE | 64.07 | 64.57 | 62.43 | 63.41 | 62.22 | 63.34 | 0.0364 | |
| Pythia-2.8B | - | Raw | 54.59 | 46.84 | 52.92 | 48.93 | 51.36 | 50.92 | 0.1184 |
| Single | Top | 68.06 | 70.84 | 60.86 | 64.93 | 62.33 | 66.13 | 0.0342 | |
| ALL | Averaging | 58.80 | 52.55 | 59.03 | 51.83 | 51.70 | 54.78 | 0.0685 | |
| ALL | MultiTask | 66.49 | 66.73 | 63.37 | 64.48 | 58.95 | 64.00 | 0.0456 | |
| ALL | MORE | 65.39 | 66.34 | 63.58 | 65.39 | 59.39 | 64.01 | 0.0366 | |
| LLaMa2-7B | - | Raw | 49.78 | 47.18 | 51.15 | 49.84 | 49.88 | 49.56 | 0.1503 |
| Single | Top | 73.08 | 74.84 | 63.58 | 67.07 | 68.65 | 69.27 | 0.0334 | |
| ALL | Averaging | 61.90 | 54.15 | 56.21 | 55.16 | 63.60 | 58.20 | 0.0391 | |
| ALL | MultiTask | 72.10 | 72.70 | 64.62 | 71.95 | 69.30 | 70.13 | 0.0570 | |
| ALL | MORE | 71.93 | 72.70 | 65.88 | 70.27 | 70.85 | 70.32 | 0.0458 | |
Training hyperparameters
All RM training batch size is set to 5 (number of preferences)*16 (batch size of each preference) = 80. For RJS experiments, we set the training batch size to 64. The max input sequence length is 512. All RMs, Alpaca-SFT, and Alpaca-RJS are finetuned with one epoch. We use optimizer AdamW (Loshchilov and Hutter, 2017) with learning rate .
Experiment platform
Our experiments are conducted on computation platform with NVIDIA A100 40G GPU * 8.
Data composition
We present the statistics of datasets in Table 3. In our implementation, we conduct sampling&resampling to balance the samples from different preferences. Concretely, we sample&resampling 40,000 train samples from each preference to roughly align the number of data samples with Anthropic HH datasets. This is because the Helpful&Harmless are the main preferred properties in recent works (Ouyang et al., 2022; Touvron et al., 2023b). Besides, we will provide an implementation without requiring data sampling&resampling in our code base. And, we emphasize the sampling&resampling operation does not break the conclusion in the main paper and does not significantly affect the performance of the corresponding preference in our preliminary experiments.
Missing experiment results
| Dataset | Num. of train samples | Num. of test samples |
| Anthropic Helpful | 43,774 | 2,352 |
| Anthropic Harmless | 42,537 | 2,312 |
| OpenAssistant Oasst1 | 18,165 | 957 |
| OpenAI Webgpt | 17,106 | 901 |
| OpenAI Summarize | 92858 | 2,000* |
| Preference | RM | Positive Outliers | Negative Outliers | |||
| Scheme | ECE | Count | Mean | Count | Mean | |
| Helpful | Top | 0.0160 | 224 | 0.866 | 70 | -0.623 |
| MultiTask | 0.0171 | 201 | 0.628 | 81 | -0.437 | |
| MORE | 0.0053 | 201 | 0.596 | 76 | -0.423 | |
| Harmless | Top | 0.0213 | 152 | 0.852 | 76 | -0.610 |
| MultiTask | 0.0183 | 146 | 0.526 | 82 | -0.411 | |
| MORE | 0.0166 | 152 | 0.523 | 72 | -0.425 | |
| Preference | RM | Positive Outliers | Negative Outliers | |||
| Scheme | ECE | Count | Mean | Count | Mean | |
| Helpful | Top | 0.0191 | 193 | 0.852 | 67 | -0.606 |
| MultiTask | 0.0147 | 198 | 0.624 | 78 | -0.417 | |
| MORE | 0.0109 | 195 | 0.640 | 81 | -0.451 | |
| Harmless | Top | 0.0057 | 132 | 0.833 | 71 | -0.608 |
| MultiTask | 0.0143 | 147 | 0.602 | 94 | -0.465 | |
| MORE | 0.0047 | 152 | 0.595 | 85 | -0.445 | |
| Preference | RM | Positive Outliers | Negative Outliers | |||
| Scheme | ECE | Count | Mean | Count | Mean | |
| Helpful | Top | 0.0081 | 223 | 1.145 | 73 | -0.784 |
| MultiTask | 0.0182 | 195 | 1.054 | 81 | -0.734 | |
| MORE | 0.0080 | 181 | 0.806 | 76 | -0.552 | |
| Harmless | Top | 0.0225 | 152 | 1.070 | 89 | -0.891 |
| MultiTask | 0.0386 | 168 | 1.206 | 77 | -0.901 | |
| MORE | 0.0206 | 156 | 1.075 | 57 | -0.821 | |
Appendix D GPT4 Evaluation
Our GPT4 evaluation aligns with the DPO (Rafailov et al., 2023). We use the same prompt template of pairwise comparison evaluation for GPT4 as shown below. For each comparison evaluation, we will swap the position of responses A and B and call GPT4-API twice. If both results are A is better, the final label will be A is better. On the contrary, the final label will be B is better. If the results are not consistent, the final label will be a tie.

### v5

@@
 However, in this pluralistic world, human preferences can be diversified due to annotators’ different tastes, which hinders the effectiveness of LLM alignment methods.
-This paper presents the first quantitative analysis of commonly used human feedback datasets to investigate the impact of diversified preferences on reward modeling.
-Our analysis reveals a correlation between the calibration performance of reward models (RMs) and the alignment performance of LLMs.
-We find that diversified preference data negatively affect the calibration performance of RMs on human-shared preferences, such as Harmless&Helpful, thereby impairing the alignment performance of LLMs.
-To address the ineffectiveness, we propose a novel Multi-Objective Reward learning method (MORE) to enhance the calibration performance of RMs on shared preferences.
-We validate our findings by experiments on three models and five human preference datasets.
-Our method significantly improves the prediction calibration of RMs, leading to better alignment of the Alpaca-7B model with Harmless&Helpful preferences.
-Furthermore, the connection between reward calibration and preference alignment performance suggests that calibration error can be adopted as a key metric for evaluating RMs.
-The open-source code and data are available at https://github.com/dunzeng/MORE.
+This paper presents the first quantitative analysis of the experimental scaling law for reward models with varying sizes, from 1.3 billion to 7 billion parameters, trained with human feedback exhibiting diverse preferences.
+Our analysis reveals that the impact of diversified human preferences depends on both model size and data size.
+Larger models with sufficient capacity mitigate the negative effects of diverse preferences, while smaller models struggle to accommodate them.
+To mitigate the impact of diverse preferences, we introduce a new metric, Expected Calibration Error (ECE), to evaluate RMs and show their obvious positive correlation with the alignment performance of LLMs.
+Furthermore, we propose a Multi-Objective Reward learning method (MORE) to enhance the calibration performance of RMs on shared preferences.
+Through experiments on four models and five human preference datasets, we find the calibration error can be adopted as a key metric for evaluating RMs and MORE can obtain superior alignment performance.
 On Diversified Preferences of Large Language Model Alignment
-Dun Zeng1,333Work done during an internship at Tencent AI Lab,111Equal contribution.
-YongDai1,111Equal contribution.
-Pengyu Cheng1,111Equal contribution.
-Longyue Wang1 Tianhao Hu1 Wanshun Chen1 Nan Du1 Zenglin Xu2,222Corresponding author: zenglin@gmail.com 1Tencent AI Lab, Shenzhen, China 2Peng Cheng Lab, Shenzhen, China
+Dun Zeng1,111Equal contribution.333Part of work was done at Tencent AI Lab YongDai2,111Equal contribution.333Part of work was done at Tencent AI Lab Pengyu Cheng1,111Equal contribution.
+Longyue Wang3,333Part of work was done at Tencent AI Lab Tianhao Hu1 Wanshun Chen1 Nan Du1 Zenglin Xu4,222Corresponding author: zenglin@gmail.com 1Tencent AI Lab, 2HiThink Research, 3Alibaba Group, 4Peng Cheng Lab
 1 Introduction
@@
 (2022); Wan et al.
+(2023); He et al.
 (2024).
@@
 To enhance the efficiency of leveraging the diversified preference datasets, inspired by multi-objective optimization methods Sener and Koltun (2018); Zeng et al.
-(2023), we regard RMs as a shared reward additionally with a customized reward drift.
+(2023c), we regard RMs as a shared reward additionally with a customized reward drift.
 The shared reward represents the shared preferences across datasets (or general human preferences) and the reward drift contains individual or domain-specific preference information (Cheng et al., 2023a).
@@
 where we adjust to minimize the partial gradient of enlarging reward drifts.
-The mitigation task in (8) can be efficiently solved by the Frank-Wolfe solver (Jaggi, 2013; Sener and Koltun, 2018; Zhou et al., 2022b; Zeng et al., 2023).
+The mitigation task in (8) can be efficiently solved by the Frank-Wolfe solver (Jaggi, 2013; Sener and Koltun, 2018; Zhou et al., 2022b; Zeng et al., 2023c).
 We provide the details of our efficient implementation in the Appendix B.
@@
 5 Experiments on Reward Modeling
-In this section, we present our experiments and quantitative analyses on reward modeling.
+In this section, we present our experiments and quantitative analyses on reward modeling. The open-source code and data are available at https://github.com/dunzeng/MORE.
 Datasets & models
@@
 Increasing the size of LLMs, reward model training on mixed diversified preference datasets can maintain reward accuracy.
-For instance, when Pythia-1.4B is used as the RM base model, the reward accuracy is lower compared to the Top accuracy achieved through single preference training on all preferences.
+For instance, when Pythia-1.4B is used as the RM base model, the reward accuracy is lower than the Top accuracy achieved through single preference training on all preferences.
 Then, when LLaMa2-7B is used as the base model, the reward accuracy on the Oasst1, Webgpt, and Summarise test sets surpasses the top accuracy achieved through single training.
@@
 Therefore, the performance of RMs typically is proportional to the size of base models (Gao et al., 2023).
-Moreover, we find the accuracy of RM is poor, revealing the preference conflicts across RM.
+Moreover, we find the accuracy of RM is low, revealing the preference conflicts across RM.
 Reward modeling on mixed diversified preferences affects calibration performance
@@
 We will further justify the point in the alignment experiments.
+We provide additional reward modeling results of LLaMa2-13B in the Appendix.
+Compared with the results of LLaMa2-7B, the reward accuracy of LLaMa2-13B is not significantly better.
+This is because the capacity of these 7B and 13B models is sufficient for fitting the datasets used.
+Notably, the ECE of the 13B model is marginally improved in the MultiTask setting, and the ECE gap between RM and RM is narrowed.
+Thus, a larger reward model can mitigate the negative impacts of mixed diverse preferences.
 Reward Model Perplexity (PPL) GPT4 Evaluation (%) Base Model Scheme Acc(%) ECE Helpful Harmless Win Tie Lose - - - - 15.48 12.71 - - - Pythia-1.4B MultiTask 64.79 0.0177 15.30 8.22 44 22 34 MORE 64.32 0.0109 12.68 8.42 Pythia-2.8B MultiTask 66.61 0.0145 16.76 8.42 45 21 34 MORE 65.87 0.0078 13.14 10.29 LLaMa2-7B MultiTask 72.40 0.0284 16.93 8.69 45 23 32 MORE 72.32 0.0143 11.97 9.96 Table 1: The RJS alignment performance with different RMs.
-The first line is the performance of Alpaca base model.
+The first line is the performance of the Alpaca base model.
 The results show that ECE further reflects the ability of RMs when the reward accuracy is close.
@@
 Then, we conduct the RJS alignment experiments with LLaMa2-7B RMs from Figure 6.
-We compare each alignment result of Alpaca-RJS models with the same Alpaca-STF model via GPT evaluation, shown in Figure 7.
+In Figure 7, we compare each alignment result of Alpaca-RJS models with the same Alpaca-STF model via GPT evaluation (the tie rates are around 15%).
 The results show that the RMs with lower ECE values work better for RJS alignments, emphasizing the importance of calibration evaluation.
@@
 In comparison, this paper focuses on training single RM on diversified datasets.
+Connections with fine-grained reward and multi-dimensional reward
+Existing research, including fine-grained reward (Wu et al., 2024) and multi-dimensional reward (Lou et al., 2024), has increasingly focused on the importance of reward diversity.
+These studies categorize the utilization of diverse reward signals into two main strategies: first, integrating multi-dimensional preferences into a single dimension for aligning large language models (LLMs); and second, decoupling preferences across dimensions (reward models) to align LLMs on each dimension.
+This body of work underscores the necessity of addressing diverse preferences.
+In contrast, our primary contribution lies in the empirical analysis of diversified preferences within a single reward model, alongside our novel evaluation of expected calibration error (ECE) on reward models.
+Additionally, since both fine-grained and multi-dimensional reward methodologies yield scalar rewards, our findings regarding ECE are relevant for assessing these approaches.
 Suggestions for reward model training
@@
 Reward accuracy is insufficient to evaluate the ability of RMs due to model capacity and data quality.
-Our work suggests the community additionally focuses on the calibration performance of RMs.
+Our work suggests that the community additionally focuses on the calibration performance of RMs.
 Besides, Increasing the diversity of preference data samples can ensure the robustness of the reward modeling process.
@@
 The MORE can enhance preference modeling pre-trained (PMP) paradigm (Askell et al., 2021) as it captures the shared preference information.
-This facilitates its use in federated learning scenarios (McMahan et al., 2017), where the data distributions are highly heterogeneous across participants.
+This facilitates its use in federated learning scenarios (McMahan et al., 2017; Zeng et al., 2023b, a), where the data distributions are highly heterogeneous across participants.
 Moreover, the RM can be easily finetuned to specific preferences (Cheng et al., 2023a).
@@
 In the practice of machine learning, most MOO methods Sener and Koltun (2018); Zeng et al.
-(2023) apply linear scalarization Barrett and Narayanan (2008) to merge multiple objectives into one, and then automatically adjust the objective coefficients to balance the conflicts among different tasks.
+(2023c) apply linear scalarization Barrett and Narayanan (2008) to merge multiple objectives into one, and then automatically adjust the objective coefficients to balance the conflicts among different tasks.
 Appendix B Detailed Discussions about MORE
@@
 We use adaptive weighting methods to reduce the reward drift across preferences and adjust the reward modeling process in the data batch-wise.
-The mitigation task in (8) can be efficiently solved by the Frank-Wolfe solver (Jaggi, 2013; Sener and Koltun, 2018; Zhou et al., 2022b; Zeng et al., 2023).
+The mitigation task in (8) can be efficiently solved by the Frank-Wolfe solver (Jaggi, 2013; Sener and Koltun, 2018; Zhou et al., 2022b; Zeng et al., 2023c).
 However, the computing cost of solving it is proportional to the size of parameters .
@@
 | ALL | MORE | 71.93 | 72.70 | 65.88 | 70.27 | 70.85 | 70.32 | 0.0458 | |
+| LLaMa2-13B | - | Raw | 50.71 | 48.47 | 50.35 | 49.87 | 49.18 | 49.71 | 0.1478 |
+| Single | Top | 75.29 | 74.82 | 65.98 | 67.07 | 71.18 | 71.46 | 0.0275 | |
+| ALL | Averaging | 56.04 | 49.38 | 56.41 | 59.32 | 45.96 | 53.42 | 0.0471 | |
+| ALL | MultiTask | 73.85 | 73.91 | 65.36 | 71.98 | 69.59 | 70.90 | 0.0561 | |
+| ALL | MORE | 73.80 | 73.22 | 64.95 | 69.51 | 70.14 | 70.32 | 0.0502 | |
 Training hyperparameters
@@
 | Scheme | ECE | Count | Mean | Count | Mean | |
-| Helpful | Top | 0.0160 | 224 | 0.866 | 70 | -0.623 |
+| Helpful | Single | 0.0160 | 224 | 0.866 | 70 | -0.623 |
 | MultiTask | 0.0171 | 201 | 0.628 | 81 | -0.437 | |
 | MORE | 0.0053 | 201 | 0.596 | 76 | -0.423 | |
-| Harmless | Top | 0.0213 | 152 | 0.852 | 76 | -0.610 |
+| Harmless | Single | 0.0213 | 152 | 0.852 | 76 | -0.610 |
 | MultiTask | 0.0183 | 146 | 0.526 | 82 | -0.411 | |
@@
 | Scheme | ECE | Count | Mean | Count | Mean | |
-| Helpful | Top | 0.0191 | 193 | 0.852 | 67 | -0.606 |
+| Helpful | Single | 0.0191 | 193 | 0.852 | 67 | -0.606 |
 | MultiTask | 0.0147 | 198 | 0.624 | 78 | -0.417 | |
 | MORE | 0.0109 | 195 | 0.640 | 81 | -0.451 | |
-| Harmless | Top | 0.0057 | 132 | 0.833 | 71 | -0.608 |
+| Harmless | Single | 0.0057 | 132 | 0.833 | 71 | -0.608 |
 | MultiTask | 0.0143 | 147 | 0.602 | 94 | -0.465 | |